Wentworth's Five and Dime Store has a cost of equity of 11.5 percent. The company has an aftertax cost of debt of 5.1 percent, and the tax rate is 39 percent. If the company's debt–equity ratio is .75, what is the weighted average cost of capital?
Answer #1

Debt/Equity = 0.75

= 75/100

Hence total value = 75+100 = 175

WACC = Weight of equity * Cost of equity + Weight of debt * cost of debt after tax

= (100/175)*11.5%+ (75/175)*5.1%

=8.76%